OpenStudy (anonymous):

Solve: 8x − 10 + 18 = 24

mathslover (mathslover):

\[\huge{8x-10+18=24}\] \[\huge{8x+8=24}\] \[\huge{8x+8-8=24-8}\] \[\huge{8x=16}\] \[\huge{x=\frac{16}{8}=2}\]

OpenStudy (anonymous):

8x = 24 + 10 - 18 8x = 34 - 18 Can you solve it further???

mathslover (mathslover):

you can see my tutorials for transposing and solving these type of equations @darkbbn

OpenStudy (anonymous):

Is it 2? :)

mathslover (mathslover):

yes

OpenStudy (anonymous):

ty got another

OpenStudy (anonymous):

Yes you are right in solving this...